OpenStudy (aaronandyson):

Ten years ago,prisha's age was the square of her sisters age.11 years hence she will be 6 years older to her sister.find their present age

Parth (parthkohli):

Very badly framed question. The square of which number is six greater than itself? \(3\). Now, they must be \(13\) and \(19\).

Parth (parthkohli):

"11 years hence she will be 6 years older to her sister.find their present age" this means that they currently differ by six years too. they will ALWAYS differ by six years, not only eleven years later.

OpenStudy (aaronandyson):

older than her sister **

Parth (parthkohli):

If you want a quadratic equation, just try this: Let the present ages of Prisha and her sister be \(x\) and \(x-6\).\[(x-6-10)^2 = x-10\]

OpenStudy (aaronandyson):

Ten years ago,Prisha's age was the square of her sisters age. Eleven years hence,she will be 6 years older than her sister. Find their present age.

Parth (parthkohli):

Yes - 19 and 13.

OpenStudy (aaronandyson):

who is 13..?

Parth (parthkohli):

Who's younger? Prisha's sister of course.

OpenStudy (aaronandyson):

-19 cant be a persons age.. O.o

Parth (parthkohli):

I never said -19. I said 19.

OpenStudy (aaronandyson):

oops... i miss read it

Parth (parthkohli):

You can check both facts in the question.

OpenStudy (aaronandyson):

I just haveone last question

Parth (parthkohli):

Ten years back, their ages were 9 and 3. And 3^2 = 9. Eleven years later, their ages will be six apart.

Parth (parthkohli):

When do you have your pre-boards?

OpenStudy (aaronandyson):

Friday ko khtm hoye hain

OpenStudy (aaronandyson):

second pre-board 4 jan 16 se hai

Parth (parthkohli):

Oh, itni jaldi? So your 10th syllabus is over?

OpenStudy (aaronandyson):

haaan

Parth (parthkohli):

I remember my pre-boards. Best exams I ever took. Still have my papers.